OpenStudy (anonymous):

the angle θ between two radii OP and OQ of a circle of radius 6 cm is increasing at the rate of 0.1 radians per minute. (a) Show that the area of sector OPQ is increasing at the rate of 1.8cm^2/min. (b) Find the rate at which the area of ∆OPQ is increasing at the instant when θ= π/4. (c) Find the value of θ for which the rate of increase of the area of the segment cut off by the cord PQ is at its maximum

OpenStudy (anonymous):

ok fine. what is the area of the sector?

OpenStudy (anonymous):

i have no idea thats all the info we get

OpenStudy (anonymous):

no i mean what is the formula for an area of a sector given the angle theta and the radius r ?

OpenStudy (anonymous):

k it is \[A=\frac{\theta}{2}r^2\]

OpenStudy (anonymous):

1/2 x θ x r^2

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

here r = 6 and \[\frac{d\theta}{dt}=0.1\]

OpenStudy (anonymous):

you want \[\frac{dA}{dt}\] which is \[\frac{dA}{dr}\times \frac{dr}{dt}\]

OpenStudy (anonymous):

oh did i mess that up. forget the latex for a second

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

you want \[\frac{dA}{dt}=\frac{dA}{d\theta}\frac{d\theta}{dt}\]

OpenStudy (anonymous):

so here \[A=18r^2\] since r = 6, r^2= 36 and 36/2=18

OpenStudy (anonymous):

damn it i must be tired. sorry

OpenStudy (anonymous):

\[A=18\theta\] for the above reasons

OpenStudy (anonymous):

oh isee

OpenStudy (anonymous):

\[\frac{dA}{d\theta}=18\]

OpenStudy (anonymous):

and \[\frac{d\theta}{dt}=.1\] because that is what you were told

OpenStudy (anonymous):

oh i get that now

OpenStudy (anonymous):

and that finally is the answer, but only to #1

OpenStudy (anonymous):

so for the next one find the area of a triangle , then diffrentiate and sub in θ= pi/4 to find the rate?

OpenStudy (anonymous):

i use 1/2 ab sin C right

OpenStudy (anonymous):

where c = θ

OpenStudy (anonymous):

area of triangle is oh yes, i was looking but of course you do not need a formua it is \[\frac{1}{2} 6^2 sin(\theta)\] yes?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

which i guess is \[18sin(\theta)\] in this case

OpenStudy (anonymous):

then it becomes 18 cos θ

OpenStudy (anonymous):

great, then take the derivative, and then multiply by .1 and then plug in \[\frac{\pi}{4}\]

OpenStudy (anonymous):

oh k cool got that one sorted now for the last one

OpenStudy (anonymous):

i am guessing \[.9\sqrt{2}\]

OpenStudy (anonymous):

now for the next one i am going to guess \[\frac{\pi}{2}\]

OpenStudy (anonymous):

\[9\sqrt{2}/10 cm^2/\min\]

OpenStudy (anonymous):

yes that is what i got too, only i wrote .9 but great!

OpenStudy (anonymous):

oh cool

OpenStudy (anonymous):

now we have the formula for the area and it is \[18\sin(\theta)\]

OpenStudy (anonymous):

do you need to find the second derivate of the area of a segment and find its maximum

OpenStudy (anonymous):

the rate of change is \[.18\cos(\theta)\]

OpenStudy (anonymous):

right you are i misread the question. you want the rate to be biggest, not the change

OpenStudy (anonymous):

so take the second derivative set = 0 and solve yes? what you said

OpenStudy (anonymous):

yay cool got it

OpenStudy (anonymous):

btw the rate of change is \[1.8\cos(\theta)\] not what i wrote earlier

OpenStudy (anonymous):

great. just about to it in your head. derivative is -sin(x) and it is 0 at 0 and pi

OpenStudy (anonymous):

not too bad this one. i looked harder at the beginning. good luck

OpenStudy (anonymous):

kk thanks heap btw impressive medal count :)